OpenStudy (anonymous):

find f '(x) of 1/square root of t. give the definition of the derivitative. f '(x)=(f(x+a)-f(x))/(h) I know you find denominator, then conjugate, but now what??

OpenStudy (dumbcow):

\[\frac{\frac{1}{\sqrt{t+h}} - \frac{1}{\sqrt{t}}}{h}\] \[= \frac{\sqrt{t} - \sqrt{t+h}}{h \sqrt{t} \sqrt{t+h}}\] like you said, multiply/divide by conjugate \[= \frac{t - (t+h)}{h \sqrt{t} \sqrt{t+h} (\sqrt{t} + \sqrt{t +h})}\] t - t = 0 leaving an h on top .... this cancels with h in denominator \[= \frac{-1}{\sqrt{t} \sqrt{t+h} (\sqrt{t} + \sqrt{t +h})}\] plug in h=0 and evaluate \[= -\frac{1}{2 t \sqrt{t}}\]

OpenStudy (anonymous):

thank you thank you!!!!